if one third of a number exceeds one forth of that number by 16 , then the number is
Answer by mananth(16949)

if one third of a number exceeds one forth of that number by 16 , then the number is
let the number be x
1/3 of the number is x/3
1/4 of the number = x/4
x/3 -x/4 =16
4x-3x / 12 =16
x= 12 *16
x= 192
the number is 192